OpenStudy (anonymous):

PLEASE HELP WILL FAN AND MEDAL factor x^2-9x+20

OpenStudy (anonymous):

@$oumya

OpenStudy ($oumya):

=x*x-9*x+20 =x*x-5*x-4*x+20 =x(x-5)-4(x-5) =(x-5) (x-4)